OpenStudy (anonymous):

Solve for t using natural logarithms. 7e^6t=10e^5t

OpenStudy (anonymous):

how?

OpenStudy (saifoo.khan):

Oh, wait.

OpenStudy (saifoo.khan):

Let me show it.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

it says to round 3 decimals to prob not 0

OpenStudy (saifoo.khan):

\[\Large e^{6t} = \frac{10}{7} e^{5t}\]

OpenStudy (saifoo.khan):

This equation thing isnt workin. just a sec.

OpenStudy (anonymous):

okk

OpenStudy (saifoo.khan):

\[\Large 6t = \ln (\frac{10}{7} e^{5t})\]

OpenStudy (anonymous):

and from there?

OpenStudy (saifoo.khan):

Thinkin.

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

wow, this is confusing.

OpenStudy (saifoo.khan):

Got it. :D

OpenStudy (anonymous):

yayyyy

OpenStudy (saifoo.khan):

\[\Large 6t = \ln (\frac{10}{7}) + \ln e^{5t}\] \[\Large 6t = \ln (\frac{10}{7}) + 5t\ln e\] \[\Large 6t = \ln (\frac{10}{7}) + 5t(1)\] \[\Large 6t = \ln (\frac{10}{7}) + 5t\] \[\Large 6t-5t = \ln (\frac{10}{7})\] \[\Large t = \ln(\frac{10}{7})\]

OpenStudy (anonymous):

sure?

OpenStudy (saifoo.khan):

100%.

OpenStudy (anonymous):

awesome! ill try it. thankss

OpenStudy (anonymous):

Wow, thanks so much :)

OpenStudy (saifoo.khan):

Welcome.